OpenStudy (anonymous):

f(x)=-x^2(x-1)(x+3) Determine the end behavior: find the power function that the graph of f resembles for large values of |x|. Find the x-and y-intercepts of the graph. Determine whether the graph crosses or touches the x-axis at each x-intercept.

ganeshie8 (ganeshie8):

look at the given polynomial f(x) whats the degree of f(x) ?

OpenStudy (anonymous):

im not given any additional information.. just was given the problem and asked to find that information..

ganeshie8 (ganeshie8):

Ahh, actually im asking you a question to see if you know how to find the "degree"

OpenStudy (anonymous):

I'm pretty sure i understand how to find the x and y intercepts.. should i treat this as a normal polynomial function and cross multiply?

OpenStudy (anonymous):

oh!! i thought would be the exponent

ganeshie8 (ganeshie8):

it is indeed a polynomial, but do not multiply yet, i think we can finish the problem with out multiplying

ganeshie8 (ganeshie8):

yes, whats the degree of f(x) ?

OpenStudy (anonymous):

i mean the degree is the exponent*

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

exponent of what

OpenStudy (anonymous):

but then again I'm thinking its just 1

OpenStudy (anonymous):

x^2?

ganeshie8 (ganeshie8):

f(x)=-x^2(x-1)(x+3) notice that there are 3 factors here, multiplying them out gives the exponent of leading term as : 2+1+1 = 4 so the degree of f(x) is 4

ganeshie8 (ganeshie8):

for part a, the power function that the graph of f resembles for large values of |x| is \(\large -x^4\)

OpenStudy (anonymous):

so as if i was to cross multiply id get an exponent of 2 for the first value and 1 and 1 on the middle and last term?

OpenStudy (anonymous):

ok...

ganeshie8 (ganeshie8):

Exactly! \(f(x)=-\color{red}{x}^2(\color{red}{x}-1)(\color{red}{x}+3)\) its easy to see that \(-\color{Red}{x}^2*\color{Red}{x}*\color{red}{x}\) gives you \(-\color{Red}{x}^{2+1+1} = -\color{red}{x}^4\)

OpenStudy (anonymous):

what does finding the degree do for me though?

ganeshie8 (ganeshie8):

good question, when you plugin large values of |x|, the function behaves almost same as the power function\(\large -x^4\)

OpenStudy (anonymous):

I've been working similar problems and haven't needed that information to find the x and y intercepts

ganeshie8 (ganeshie8):

that degree thing helps in answering part a of question : `Determine the end behavior: find the power function that the graph of f resembles for large values of |x|.`

OpenStudy (anonymous):

wow... ok

ganeshie8 (ganeshie8):

you can see here, how closely the graph of f(x) matches with the power function \(-x^4\) https://www.desmos.com/calculator/vj2xhutn6z

OpenStudy (anonymous):

i havent been shown how to find the solution to these questions except for finding the x and y intercepts, domain/range, but not the power function.

OpenStudy (anonymous):

ok yes!! i see that now.

OpenStudy (anonymous):

ok so finding the degree will give me the power function..

OpenStudy (anonymous):

Sorry lot of questions... i think I'm just having a hard time understanding what I'm being asked to find.

ganeshie8 (ganeshie8):

good, lets find the x intercepts

OpenStudy (anonymous):

ok so from what I've understood, to find the x-intercept id want to find the value of the equation by f(x)=0, make the equation equal to 0, and finding the value of x.. thats why i assumed id cross multiply..

OpenStudy (anonymous):

but now I'm not seeing how that would work.

ganeshie8 (ganeshie8):

thats right! to find x intercepts, you want to solve \[\large -x^2(x-1)(x+3)=0\]

OpenStudy (anonymous):

is this problem similar to the problem f(x)={3+x If x<0 // x^2 if x> not equal to 0....

ganeshie8 (ganeshie8):

use zero product property to get : \(-x^2=0\) or \(x-1= 0\) or \(x+3=0\)

ganeshie8 (ganeshie8):

that gives you x intercepts : \(0, 1, -3\)

OpenStudy (anonymous):

awesome!! ok so thats what i was thinking..

OpenStudy (anonymous):

then id have to find the y intercept?

ganeshie8 (ganeshie8):

yes, simply put \(x=0\) in f(x) to get the y intercepts

ganeshie8 (ganeshie8):

f(x)=-x^2(x-1)(x+3) f(0) = ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

wouldn't i substitue all x values with 0?

ganeshie8 (ganeshie8):

yes, but -3 is wrong

ganeshie8 (ganeshie8):

f(x)=-x^2(x-1)(x+3) f(0) = -0^2(0-1)(0+3) = ?

OpenStudy (anonymous):

+3

ganeshie8 (ganeshie8):

no, try again

ganeshie8 (ganeshie8):

what do you get when u multiply something by 0 ?

OpenStudy (anonymous):

0

ganeshie8 (ganeshie8):

so..

OpenStudy (anonymous):

so the answer is 0

OpenStudy (anonymous):

cuz im retarded and looked at the problem as if the 0 in front of the parentheses was invisible lol...

ganeshie8 (ganeshie8):

haha happens :)

ganeshie8 (ganeshie8):

so the y intercept is 0

ganeshie8 (ganeshie8):

how do we answer the last part `Determine whether the graph crosses or touches the x-axis at each x-intercept.`

OpenStudy (anonymous):

plot the points and read the graph and how it falls

ganeshie8 (ganeshie8):

look at the graph https://www.desmos.com/calculator/bdlowvla6f

ganeshie8 (ganeshie8):

notice that the graph is crossing x axis at x=-3 and x=1 but it is just touching x axis at x=0

OpenStudy (anonymous):

ok so it touches the x axis at 0,1,-3 but only crosses at 1 and -3

OpenStudy (anonymous):

ooh you answered before i could finish typing.. but that is what i saw

ganeshie8 (ganeshie8):

very good!

OpenStudy (anonymous):

but then how do know the shape of the graph just by the points.. how do we know how far up it goes and where it lies on the rest of the graph

ganeshie8 (ganeshie8):

the graph goes forever on both ends, it sinks down :